# Federated Variational Inference for Bayesian Mixture Models

### v1

@@ Abstract
We present a federated learning approach for Bayesian model-based clustering of large-scale binary and categorical datasets. We introduce a principled ‘divide and conquer’ inference procedure using variational inference with local merge and delete moves within batches of the data in parallel, followed by ‘global’ merge moves across batches to find global clustering structures. We show that these merge moves require only summaries of the data in each batch, enabling federated learning across local nodes without requiring the full dataset to be shared. Empirical results on simulated and benchmark datasets demonstrate that our method performs well in comparison to existing clustering algorithms. We validate the practical utility of the method by applying it to large scale electronic health record (EHR) data.
Machine Learning, ICML
@@ 1 Introduction
Federated learning (FL) refers to the setting in which a network of clients collaboratively train a model, while ensuring that the training data are kept local (Kairouz et al., 2021). FL was introduced as a way to enable learning when datasets are large and/or privacy sensitive (McMahan et al., 2017). There have been numerous developments in supervised FL, but methods for unsupervised FL remain limited (Tian et al., 2024). While a number of federated clustering algorithms have been proposed (e.g. federated $k$-means: Kumar et al., 2020; Liu et al., 2020; Garst & Reinders, 2025), and Bayesian FL (BFL) is an active area of research (see Cao et al., 2023, for a survey), we are not aware of previous studies into FL for Bayesian model-based clustering, which we consider here.
@@
Mixture models provide a probabilistic approach for clustering by assuming data are generated from a mixture of underlying probability distributions. Bayesian mixture models enable the uncertainty in cluster allocations and, potentially, the number of clusters to be modelled. These include Bayesian finite mixture models (e.g. Diebolt & Robert, 1994; Richardson & Green, 1997; Rousseau & Mengersen, 2011), as well as Bayesian nonparametric methods, such as Dirichlet process (DP) mixture models (Ferguson, 1973; Escobar & West, 1995). Markov Chain Monte Carlo (MCMC), particularly Gibbs sampling (Neal, 2000), is often used to draw samples from the intractable posterior (e.g. Robert & Casella, 2004; Walker, 2007). Despite efforts to improve scalability, MCMC algorithms are often too computationally intensive for even moderately large datasets, and/or may experience mixing issues, preventing the sampler from reaching convergence (Celeux et al., 2000).
To address the growth of ‘massive’ datasets, a range of methods have been proposed to scale MCMC algorithms to large sample sizes (e.g Huang & Gelman, 2005). These include distributed computing approaches in which the computational and memory requirements of an algorithm are split, with the algorithm running in parallel on multiple nodes (Zhu et al., 2017). An important example of this for Bayesian inference is consensus Monte Carlo, where a Monte Carlo algorithm is run on partitions of the data and samples are then aggregated across partitions to enable ‘global’ inferences to be drawn (Scott et al., 2016). Various methods have been proposed to aggregate the subset posteriors. Informally, the challenge for clustering models is how to match clusters/cluster labels across data partitions. The SNOB (Zuanetti et al., 2018) and SIGN (Ni et al., 2019) algorithms have been proposed to address this.
Distributed variational inference (VI) algorithms have also been proposed to obtain an approximation of the posterior distribution, including variational consensus Monte Carlo (VCMC) (Rabinovich et al., 2015). As noted in Rabinovich et al. (2015), models with unidentifiable latent variables, such as component allocation labels in mixture models, present additional complexity for such approaches, and the authors suggested a heuristic approach using the Hungarian algorithm (Kuhn, 1955) to align cluster centres across data partitions. Our proposed method instead utilises the variational objective function to allow us to combine results from data partitions in a statistically principled manner.
@@
While VI methods tend to have better run-time performance than MCMC algorithms (e.g. Blei et al., 2017), they can also be sensitive to initialisation due to local optima in the objective function. To address these issues when fitting mixture models, ‘merge’ and ‘delete’ moves may be incorporated into the variational algorithm to dynamically combine or remove unnecessary clusters. Such moves, in addition to ‘split’ moves to create new clusters, have been seen to improve inference in expectation-maximisation (EM) algorithms (Ueda et al., 2000) and MCMC samplers (Jain & Neal, 2004; Dahl & Newcomb, 2021) for mixture models, and have also been employed less commonly in VI algorithms for DP mixtures, including Gaussian DP mixture models (Hughes et al., 2015; Lin, 2013), and the hierarchical DP (Hughes & Sudderth, 2013).
Here we propose a distributed algorithm that performs inference for Bayesian mixture models using VI with local merge and delete moves within batches of the data across a network of nodes in parallel, followed by ‘global’ merge moves to aggregate clusters across batches. We show that these ‘global’ moves require only summaries of the data from each node, enabling federated learning of the model without requiring the full dataset to be shared. We anticipate that the approach may be useful for the analysis of privacy sensitive datasets, and provide an example application clustering individual-level electronic health records (EHRs).
We focus on applications to categorical, particularly binary (2-category) datasets, but note that our approach can be straightforwardly adapted to mixtures of Gaussians and other members of the exponential family. In Section 2, we provide an overview of variational Bayesian mixture models, and in Section 3 we briefly introduce merge/delete moves. Section 4 describes how we can employ these moves to enable scaling to large datasets and federated learning. Sections 5 and 6 present results from both simulated and real-world data, including a British EHR dataset, and we discuss the results and take a brief look at future directions for the model in Section 7. We consider in the Appendix how our model can be extended to tackle variable selection.
@@ 2.1 Bayesian Finite Mixture Models
$p(X|\boldsymbol{\pi},\boldsymbol{\phi})=\sum_{k=1}^{K}\pi_{k}f(X|\phi_{k}),$ (1)
where component densities $f(X|\phi_{k})$ are usually from the same parametric family but with different parameters associated with each component. In the limit as $K$ goes to infinity, this is equivalent to the Dirichlet process (DP) mixture model (Maceachern, 1994; Escobar & West, 1995). Our model can be seen as a finite truncation of the DP mixture model to allow for simpler inference, setting $K$ to be larger than the number of clusters we expect. Each component is modelled as a categorical distribution across $P$ covariates with parameters $\phi_{kj}=[\phi_{kj1},...,\phi_{kjL_{j}}]$, where $\phi_{kjl}$ represents the probability that variable $j$ takes value $l$ in component $k$. $L_{j}$ is the number of categories for variable $j$. The mixture weight for the $k$-th component is denoted by $\pi_{k}$, satisfying $\sum_{k=1}^{K}\pi_{k}=1$ and $\pi_{k}\geq 0$.
@@
where all Dirichlet priors are symmetric. It has been shown theoretically that if $K$ exceeds the true number of clusters, then, under certain assumptions, setting $\alpha_{0}<1$ in Equation (3) allows the posterior to asymptotically converge to the correct number of clusters as the number of observations goes to infinity (Rousseau & Mengersen, 2011; van Havre et al., 2015; Malsiner-Walli et al., 2014). Further details of hyperparameter settings are provided in the Appendix. The model can be extended to simultaneously tackle variable selection, as in (Law et al., 2004) and (Tadesse et al., 2005), which is explored in the Appendix.
@@ 2.2 Mean-Field Variational Inference (VI)
We employ a variational inference (VI) approach to seek an approximate distribution $q$ to the posterior. We optimise the Evidence Lower Bound (ELBO) given below, $\mathcal{L}(q)$, with respect to $q(\theta)$, where $\theta$ is a collection of all parameters in the model.
[…]
This is equivalent to minimising the Kullback-Leibler (KL) divergence between $q(\theta)$ and the true posterior, $p(\theta|X)$. We constrain $q$ to be a mean-field approximation which can be fully factorised as: $q(\theta)=q_{Z}(Z)q_{\pi}(\pi)q_{\Phi}(\Phi)$. We use a standard variational EM algorithm to optimise $\mathcal{L}(q)$ (see, for example, Bishop, 2006), as described below.
@@
$\alpha^{\ast}_{k}$ $=\alpha_{0}+\sum_{n=1}^{N}r_{nk}$ (10)
$\epsilon^{\ast}_{kjl}$ $=\epsilon_{jl}+\sum_{n=1}^{N}\mathbb{I}(x_{nj}=l)r_{nk}$ (11)
Expectations throughout are taken with respect to variational distributions. Between each ‘E’ and ‘M’ step, we calculate the value of the ELBO and use this to monitor convergence. All values are initialised using k-modes (Chaturvedi et al., 2001), a method analogous to k-means for categorical data.
@@ 3 Variational Merge/Delete Moves
As in eg. (Hughes et al., 2015), we introduce merge and delete moves which can be integrated into the VI algorithm. Merge moves refer to moves that combine observations from two clusters into one unified cluster; delete moves refer to moves to delete unnecessary extra clusters in the model. Both moves enable redundant clusters to be efficiently removed to help the model to escape local optima, and are designed in a principled manner to align with the ELBO variational objective function.
When moves are proposed, parameters in the VI framework are updated through initially updating responsibilities, $r_{nk}$, which are directly associated with the cluster assignment of observations, and then following through with classical variational EM steps. Moves are only accepted if they improve or maintain the ELBO, otherwise the original configuration is maintained, ensuring that the method remains consistent with the underlying VI framework.
Variational implementations of these moves in Dirichlet process mixture models, in addition to split moves, have been shown to outperform standard variational Bayes and related algorithms (Hughes et al., 2015; Yang et al., 2019). However, little work has been published regarding variational categorical mixture models. In particular, split moves are not relevant in the overfitted finite mixture model setting. We provide clear, practical strategies for the implementation of these moves in the categorical setting in Appendix E.
In particular, we consider how candidate clusters for merge and delete moves can be chosen to enhance efficiency in categorical mixture models. Comparing ELBOs between all possible merge/delete proposals would result in an unnecessarily high computational cost. We demonstrate through simulations in Appendix J.4 that using fast-to-calculate heuristics such as correlation between cluster parameters (when proposing to merge) allows for enhanced efficiency and a higher percentage of accepted merge moves compared to randomly selecting candidate clusters. This aspect has rarely been explored in prior research.
3.1 Computational Considerations
We name the variational algorithm including merge/delete moves ‘MerDel’ and introduce a parameter, ‘laps’, representing how many variational EM cycles we go through before performing a merge and a delete move.
Both merge and delete moves include some variational E and M steps to reassign observations to new clusters as described in Appendix E. These merge/delete moves are therefore relatively costly if the move is ultimately rejected, as opposed to doing purely variational steps which are guaranteed to improve the ELBO. As the model approaches convergence, more variational moves are needed to fine-tune cluster assignments while merge/delete proposals are rejected and increase overall computation. There is subsequently a trade-off between proposing more merge/delete moves to allow for bigger, more effective moves earlier on in MerDel versus reducing the amount of ‘wasted’ moves as the model approaches convergence, which we examine further in simulations. One method to mitigate this in future could be to dynamically change the ‘laps’ parameter as merges/deletes are more frequently rejected.
@@ 3.2 Parallelising MerDel
VI requires fewer sweeps over the dataset to perform inference compared to MCMC samplers, and is feasible for datasets with thousands of observations. This allows variational mixture models to be more computationally efficient and converge quickly, especially with the addition of merge/delete moves allowing models to avoid getting stuck in local optima. Despite these improvements, VI still reaches computational bottlenecks when $N$ is particularly large; we see in simulations later that MerDel is infeasible as we reach datasets with $N$ of order $10^{5}$. Repeatedly updating $N\times K$ responsibilities every E step, and subsequently summing these over $N$ to update cluster-specific parameters every M step is computationally costly once $N$ is large, even if there are 100s of iterations compared to the 1000s seen in MCMC.
All individual components of the original algorithm can be parallelised using standard methods. For example, in the E step, calculation of responsibilities $r_{nk}$ can be parallelised over $n$. Summations in the M step can also be parallelised. We provide an implementation of such a parallelised VI as part of this manuscript.
However, the overall algorithm is not embarrassingly parallel; for example, $r_{nk}$ values for all $N$ observations are required to update cluster-specific parameters for a given cluster $k$. Prior literature has consistently highlighted the limited gains achieved by such partial parallelisation for iterative methods such as VI. This is due to the significant computational overhead incurred when feeding intermediate results to and from one ‘master’ core - a globally shared computation unit - at regular, frequent intervals (Nallapati et al., 2007; Neiswanger et al., 2015). Nallapati et al. in particular saw that when parallelising variational E and M steps in Latent Dirichlet Allocation, the gain in efficiency was not significant due to the read-conflict between various threads, and also raised concerns about memory constraints as the whole dataset was stored in memory (2007).

### v2

@@ Abstract
We present a one-shot, unsupervised federated learning approach for Bayesian model-based clustering of large-scale binary and categorical datasets, motivated by the need to identify patient clusters in privacy-sensitive electronic health record (EHR) data. We introduce a principled ‘divide-and-conquer’ inference procedure using variational inference with local merge and delete moves within batches of the data in parallel, followed by ‘global’ merge moves across batches to find global clustering structures. We show that these merge moves require only summaries of the data in each batch, enabling federated learning across local nodes without requiring the full dataset to be shared. Empirical results on simulated and benchmark datasets demonstrate that our method performs well relative to comparator clustering algorithms. We validate the practical utility of the method by applying it to a large-scale British primary care EHR dataset to identify clusters of individuals with common patterns of co-occurring conditions (multimorbidity).
keywords:
Federated learning, Bayesian inference, cluster analysis
Data and Code Availability
Code is available here, which includes code for generation of all simulated datasets and random seeds used in this paper.
The MNIST dataset used is publicly available (http://yann.lecun.com/exdb/mnist/), and details and code for preprocessing are available in the above repository. The THIN EHR dataset used is not publicly available, but we provide details regarding preprocessing in the Appendix.
Institutional Review Board (IRB)
Collection of data in THIN was approved by the NHS South East Multi-Centre Research Ethics Committee (MREC). Approval to conduct this analysis was obtained from the Scientific Review Committee (reference number: 21SRC055).
@@ 1 Introduction
Federated learning (FL) refers to the setting in which a network of clients collaboratively train a model while keeping training data local (Kairouz2021), introduced for privacy-sensitive datasets (McMahan2017) common in healthcare. While supervised FL has seen many developments, unsupervised FL methods remain limited (Tian2024). With the primary motivation of clustering electronic health record (EHR) data across multiple hospitals, we address: (i) preserving data privacy while learning global clustering structure, (ii) enabling scalability to large patient record datasets, and (iii) maintaining clustering accuracy in distributed environments. While various federated clustering algorithms and Bayesian FL methods exist (IJCAI2023), there is limited work on FL for unsupervised Bayesian model-based clustering. Our method implements a one-shot FL scheme where clients perform local computation and share summary statistics with a central server in a single round, enabling asynchronous participation ideal for healthcare institutions. By combining clustering information across sites, we identify population-level disease clusters transcending institutional boundaries, enabling characterization of multimorbidity patterns across diverse populations impossible with site-specific clustering alone.
@@
Mixture models provide a probabilistic approach for clustering by assuming data are generated from a mixture of underlying probability distributions. Bayesian mixture models enable the uncertainty in cluster allocations and, potentially, the number of clusters to be modelled. Markov Chain Monte Carlo (MCMC), particularly Gibbs sampling (Neal2000), is often used to draw samples from the intractable posterior (e.g. Robert2004; Walker2007). Despite efforts to improve scalability, MCMC algorithms are often too computationally intensive for even moderately large datasets, and/or may experience mixing issues, preventing the sampler from reaching convergence (Celeux2000).
To address the growth of ‘massive’ datasets, a range of methods have been proposed to scale algorithms to large sample sizes (e.g Huang2005). These include distributed computing approaches in which the computational and memory requirements of an algorithm are split, with the algorithm running in parallel on multiple nodes (Zhu2017). However, models with unidentifiable latent variables present additional complexity for such approaches. Furthermore, few distributed computing approaches are in the spirit of FL, where data heterogeneity, data privacy, asynchronous client participation and node dropout are challenges to address.
@@
While variational inference (VI) methods tend to have better run-time performance than MCMC algorithms (e.g. Blei2017), they can also be sensitive to initialisation due to local optima in the objective function. To address these issues when fitting mixture models, ‘merge’ and ‘delete’ moves may be incorporated into the variational algorithm to dynamically combine or remove unnecessary clusters. Such moves, in addition to ‘split’ moves to create new clusters, have improved inference in expectation-maximisation (EM) algorithms (ueda_smem_2000) and MCMC samplers (jain_split-merge_2004; Dahl2021) for mixture models, and have also been employed less commonly in VI algorithms for Dirichlet process (DP) mixtures, including DP mixtures of Gaussians (pmlr-v38-hughes15; NIPS2013_8c19f571).
Here we propose a ‘one-shot’ distributed algorithm that performs inference for Bayesian mixture models using VI with local merge and delete moves within batches of the data across a network of nodes in parallel. This is followed by ‘global’ merge moves to aggregate clusters across batches, utilising the variational objective function to combine results from data partitions in a statistically principled manner.
We focus on applications to categorical, particularly binary (2-category), datasets, as is common in EHR datasets such as the one studied later in the manuscript. Binary data analysis has a long history in the context of characterising multimorbidity across populations (see e.g. Cornell2009-wb; Ng2018-sl; Nichols2022-hq) but there has been limited research into clustering methodology for binary and categorical data. By applying our method to large-scale EHR data, we demonstrate how global clustering structures (in this case, population-level multimorbidity patterns) can be identified without compromising data privacy. Our approach can be adapted to other members of the exponential family, and we consider in the Appendix how our model can be extended to tackle variable selection.
2 Related Work
A central challenge in scaling and distributing Bayesian inference is maintaining statistical accuracy while distributing computation. One could consider parallelising steps within VI to enhance scalability, and we look at this in Section 4.1. In the MCMC setting, an important example of distributed inference is consensus Monte Carlo (Scott2016), where Monte Carlo sampling is performed independently on partitions of the data, and the results are then aggregated to approximate a global posterior. While effective for certain models, these approaches typically struggle with models involving unidentifiable latent variables. Informally, the challenge for clustering models is how to match clusters/cluster labels across data partitions. pmlr-v37-gea15 considers a distributed inference algorithm for the DP and HDP mixture models with a variant of the DP slice sampler, but requires frequent communication between servers. The SNOB (Zuanetti2018) and SIGN (Ni2019) algorithms also perform inference for distributed mixture models, where MCMC samplers are run locally on partitions of the data, before local clusters are combined into global clusters with another sampler, allowing for asynchronous operation. These models are conceptually closest to our proposed model, although they do not consider the federated learning (FL) setting.
Distributed VI algorithms have also been proposed to obtain an approximation of the posterior distribution for mixture models, including variational consensus Monte Carlo (VCMC) (NIPS2015_e94550c9), where the authors suggested a heuristic approach using the Hungarian algorithm (Kuhn1955) to align cluster centres across data partitions. Our proposed method instead utilises the variational objective function to combine results from data partitions in a statistically principled manner. NIPS2015_38af8613 proposes a streaming, distributed VI method for the DP mixture model, but does not address data privacy. pmlr-v89-zhang19c extends stochastic VI (JMLR:v14:hoffman13a) to a distributed setting for the inference of mixture models, but relies on frequent parameter exchange between processors, despite efforts to make the algorithm asynchronous.
The majority of work on Bayesian inference in FL targets supervised tasks (see IJCAI2023, for a survey). These approaches have been extended to tackle e.g. Bayesian neural networks for local classification or regression tasks (pmlr-v222-bhatt24a), and Bayesian nonparametric methods for dynamic learning of local models (NEURIPS2021_46d0671d). Some works explore clustering clients in FL to tackle data heterogeneity in supervised settings (Sattler2021; Briggs2020). This is popular in personalised FL, concerning the learning of individual local distributions for clients in a federated network. For example, in Zhang2023, clients in the same cluster share a prior, and pmlr-v202-wu23z represents clients with mixtures of shared-parameter distributions where clients have different mixture weights. These methods aim to improve performance on downstream prediction tasks, but do not cluster data itself, differing fundamentally from our objective.
Far fewer works have explored unsupervised FL. Most adopt heuristic approaches like federated $k$-means (Kumar2020; Liu2020; Garst2025; pmlr-v139-dennis21a) or fuzzy clustering (Stallman2022; Pedrycz2022) lacking probabilistic foundations, while FedEM (NEURIPS2021_f740c8d9) provides probabilistic modeling but requires frequent communication rounds.
Overall, our work fills a gap at the intersection of unsupervised FL, Bayesian clustering, and scalable distributed inference. To the best of our knowledge, we are the first to propose a communication-efficient, one-shot federated algorithm for Bayesian model-based clustering of categorical and binary data.
@@ 3.1 Bayesian Finite Mixture Models
$p(X|\boldsymbol{\pi},\boldsymbol{\phi})=\prod_{n=1}^{N}p(x_{n}|\boldsymbol{\pi},\boldsymbol{\phi}),$ (1)
$\mbox{where}\hskip 17.00024ptp(x_{n}|\boldsymbol{\pi},\boldsymbol{\phi})=\sum_{k=1}^{K}\pi_{k}f(x_{n}|\phi_{k}),$ (2)
where component densities $f(x_{n}|\phi_{k})$ are usually from the same parametric family but with different parameters associated with each component. Under certain assumptions, the Dirichlet process (DP) mixture model may be derived by considering the limit as $K$ goes to infinity (Maceachern1994; Escobar1995; Neal2000). We consider the so-called ‘overfitted mixtures’ setting in which $K$ is set to be larger than the number of clusters we expect, which enables the number of clusters to be inferred (Rousseau2011). Each component is modelled as a categorical distribution across $P$ covariates with parameters $\phi_{kj}=[\phi_{kj1},...,\phi_{kjL_{j}}]$, where $\phi_{kjl}$ represents the probability that variable $j$ takes value $l$ in component $k$. $L_{j}$ is the number of categories for variable $j$. The mixture weight for the $k$-th component is denoted by $\pi_{k}$, satisfying $\sum_{k=1}^{K}\pi_{k}=1$ and $\pi_{k}\geq 0$.
@@
where all Dirichlet priors are symmetric. It has been shown theoretically that if $K$ exceeds the true number of clusters, then, under certain assumptions, setting $\alpha_{0}<1$ in Equation (4) allows the posterior to asymptotically converge to the correct number of clusters as the number of observations goes to infinity (Rousseau2011; vanHavre2015; MalsinerWalli2014). Further details of hyperparameter settings are provided in the Appendix. The model can be extended to simultaneously tackle variable selection, as in Law2004 and Tadesse2005, which is explored in the Appendix. This accounts for noisy features not contributing to the clustering structure, which can help address data corruption in real-world data.
@@ 3.2 Mean-Field Variational Inference (VI)
We employ a variational inference (VI) approach to seek an approximate distribution $q$ close to the true posterior. We optimise the Evidence Lower Bound (ELBO) below, $\mathcal{L}(q)$, with respect to $q(\theta)$, where $\theta$ is a collection of all parameters in the model.
$\mathcal{L}(q)=\int q(\theta)\ln\left(\frac{p(X,\theta)}{q(\theta)}\right)d\theta$ (6)
This is equivalent to minimising the Kullback-Leibler (KL) divergence between $q(\theta)$ and the true posterior, $p(\theta|X)$. We constrain $q$ to be a mean-field approximation which can be fully factorised as: $q(\theta)=q_{Z}(Z)q_{\pi}(\pi)q_{\Phi}(\Phi)$. We optimise $\mathcal{L}(q)$ using a standard variational EM algorithm, also known as Coordinate Ascent Variational Inference (see, for example, Bishop2006), as described below.
@@
$q^{\ast}(\pi)$ $=\mbox{Dirichlet}(\alpha^{\ast}_{1},...,\alpha^{\ast}_{K}),$ (9)
$\hskip 17.00024ptq^{\ast}(\phi)$ $=\prod_{k=1}^{K}\prod_{j=1}^{P}\mbox{Dirichlet}(\epsilon^{\ast}_{kj1},...,\epsilon^{\ast}_{kjL_{j}})$ (10)
@@
$\alpha^{\ast}_{k}$ $=\alpha_{0}+\sum_{n=1}^{N}r_{nk},\hskip 17.00024pt\epsilon^{\ast}_{kjl}=\epsilon_{jl}+\sum_{n=1}^{N}\mathbb{I}(x_{nj}=l)r_{nk}$ (11)
Expectations throughout are taken with respect to variational distributions, with ELBO monitored for convergence. All values are initialised using k-modes (Chaturvedi2001). The algorithm terminates when the ELBO is within a certain tolerance for 3 iterations in a row.
@@ 4 Variational Merge/Delete Moves
As in e.g. pmlr-v38-hughes15, we introduce merge and delete moves which can be integrated into the VI algorithm. Merge moves combine observations from two clusters into one unified cluster; delete moves remove unnecessary extra clusters. Both moves effectively eliminate redundant clusters to help the model to escape local optima, and are designed in a principled manner to align with the ELBO variational objective function. When moves are proposed, parameters in the VI framework are updated through initially updating responsibilities, $r_{nk}$ which are directly associated with cluster assignment. For example, to merge clusters $k_{1}$ and $k_{2}$, we set $r_{nk^{\ast}}=r_{nk_{1}}+r_{nk_{2}}$ for a new unified cluster $k^{\ast}$, and perform a variational M step to update $\pi,\phi$. We perform additional E and M steps to allow observation reassignment, and accept moves only if they improve the ELBO, otherwise we maintain the original configuration.
Variational implementations of these moves in Dirichlet process mixture models, in addition to split moves, have been shown to outperform standard variational Bayes and related algorithms (pmlr-v38-hughes15; Yang2019). However, little work has been published regarding variational categorical mixture models. Clear, practical strategies for the implementation of these moves in the categorical setting are provided in Appendix C. One of our key contributions lies in developing efficient strategies specifically for categorical mixture models to select candidate clusters to merge/delete. Comparing ELBOs between all possible merge/delete proposals would result in a high computational cost. We demonstrate in Appendix H.4 that using fast-to-calculate heuristics such as correlation between cluster parameters or divergence between variational distributions when proposing to merge allows for enhanced efficiency. This aspect has rarely been explored.
We name the variational algorithm including merge/delete moves ‘MerDel’ and introduce a parameter, ‘laps’, representing how many variational EM cycles we go through before performing a merge and a delete move. Since merge and delete moves include additional E and M steps (see Appendix C), they are costly compared to standard variational steps which guarantee ELBO improvement. Subsequently, there is a trade-off between proposing more merge/delete moves to allow for bigger, more effective moves earlier on in MerDel versus reducing wasted computation as the model approaches convergence and moves are rejected, which we examine in simulations. One method to mitigate this in future could be to dynamically change the ‘laps’ parameter.
@@ 4.1 Parallelising MerDel
VI requires fewer sweeps over the dataset to perform inference compared to MCMC samplers, and is feasible for datasets with thousands of observations. This allows variational mixture models to be more computationally efficient, especially with merge/delete moves. However, VI still reaches computational bottlenecks when $N$ is very large; we see in simulations later that MerDel is infeasible as we reach datasets with $N$ of order $10^{5}$. Repeatedly updating $N\times K$ responsibilities every E step, and subsequently summing these over $N$ every M step is costly.
Individual components of the original algorithm can be parallelised using standard methods e.g., in the E step, calculating $r_{nk}$ can be parallelised over $n$. We provide an implementation of such a parallelised VI as part of this manuscript. However, the overall algorithm is not embarrassingly parallel; for example, $r_{nk}$ values for all $N$ observations are required to update cluster-specific parameters for a given cluster $k$. Prior literature has consistently highlighted the limited gains achieved by such partial parallelisation for iterative methods such as VI. This is due to the significant computational overhead incurred when feeding intermediate results to and from one ‘master’ core - a globally shared computation unit - at frequent intervals (Nallapati2007; neiswanger2015embarrassinglyparallelvariationalinference).